In a test of Upper H 0​: mu=100 against Upper H Subscript a​: mu not equals​100, the sample data yielded the test statistic z=2.24. Find the Upper P​-value for the test.
Answer #1

Solution:

Given that,

This is two tailed test.

P ( Z > 2.24 )

Using standard normal table

1- P ( Z < 2.24) = 1-0.9875 =0.0125

P value = 0.0125